U Question 9 Write the inverse function of w(x) = -Væ– 3+1. Dont forget to switch to inverse notation at the end.
Answer #1

w(x) = -x-3 +! let y = -√2-3 tl yol 22-3 1-y 3 2 T -/x »-(4-1) = 50-3 VX-3 (1-y) > (4-43°+3 (1-4)2+3 f(be) = (1-6)2+3 - 26-3